OpenStudy (madmerc):

Write the equation of the line that passes through (1, 3) and has a slope of 2 in point-slope form.

OpenStudy (jennithemeani):

First you have to find the b, do you know how?

OpenStudy (jennithemeani):

oops nope nvm you do not need the b for this one

OpenStudy (jennithemeani):

\[y-y^1= m (x-x^1)\] this is point slope form so just fill it out

OpenStudy (jennithemeani):

y-3=3(x-1)

OpenStudy (madmerc):

okay but what do I do with that?

OpenStudy (jennithemeani):

Nothing, that is the answer in point slope form. You don't have to do anything else

OpenStudy (madmerc):

these are my choices: y – 1 = 2(x – 3) y – 3 = 2(x – 1) x – 1 = 2(y – 3) x – 3 = 2(y – 1)

OpenStudy (jennithemeani):

It the second one. Sorry I accidentally made the m 3 instead of 2

OpenStudy (madmerc):

that's okay! thanks for your help!